You are given the following total costs for a production facility for digital watches:

Number of units
Produced ('000)
0 10 20 30 40
Total Cost ($'000) 400    1,400 2,000 2,500 3,500


Assume the production facility produces where average variable costs are minimized. If the price is $80 per watch, what is your gain or loss? (If a loss, enter it as a negative number. Don't enter this as 'thousands of dollars' -- take into account that costs and output units are in thousands. For example, if the answer is $800,000, enter that as 800000, not as 800.)
Answer #1
Q (in thousand) TC (in thousand) TFC (in thousand) TVC=TC-TFC (in thousand) AVC=TVC/Q
0 400 400 0 -
10 1400 400 1000 100
20 2000 400 1600 80
30 2500 400 2100 70
40 3500 400 3100 77.5

Production facility produces where AVC are minimized. This means Q=30 units. If P=$80, then Total revenue = (30,000)(80)= $24,00,000 and total cost = $2500,000. This implies that there is a loss of $(2400,000-2500,000)= -100000.